OpenStudy (anonymous):

PLEASE HELP ME!!! I DONT GET IT PLEASE >_< how many times does the graph of y=3x^2+4x-1 intersect the x-axis

OpenStudy (anonymous):

find the x intercepts

OpenStudy (anonymous):

make y=0

OpenStudy (anonymous):

O yea and also these are the A)three B)two C)one D)none

OpenStudy (anonymous):

how can I make them y=0

OpenStudy (anonymous):

well proabola usually has 2

OpenStudy (anonymous):

parabola*

OpenStudy (campbell_st):

use the discriminant which is \[b^2 - 4ac\] when the general form of the quadratic is ax^2 + bx + c in your question a = 3, b = 4 and c = -1 then 4^2 - 4 x 3 x (-1) = 28 since discriminant is > 0 then it cuts the x axis twice

OpenStudy (anonymous):

u can also try factoring

OpenStudy (thomas5267):

Solve this. \[ y=3x^2+4x-1\;\text{when}\;y=0 \]

OpenStudy (anonymous):

wut cambell said would be correct use the discriminant of the quadratic formula b^2-4ac

OpenStudy (anonymous):

ok I havent learned discriminant of the quadratic formula

OpenStudy (campbell_st):

the conditions of the deiscriminant are \[b^2 - 4ac > 0\] there are 2 unequal roots. If the discriminant is a square number the roots are unequal and rational \[b^2 - 4ac = 0\] one repeated root \[b^2 - 4ac < 0\] no real roots

OpenStudy (anonymous):

u should've learned how to factor atleast therefore when finding the x coordinates it means that that point is on the x axis which means that the f(x) passes through the x axis

OpenStudy (campbell_st):

if you haven't learnt the discriminant then use the general quadratic formula to find the roots \[x = (-4\pm \sqrt{4^2 - 4 \times3\times-1})/(2\times6)\]

OpenStudy (anonymous):

ok that works i know that one general quadratic formula to find the roots ill do that

OpenStudy (anonymous):

I will also try and factor

OpenStudy (campbell_st):

factoring won't work